Real GDP in the United States was $17,688.89 billion in 2016 and $18,108.082 billion in 2017....

  1. Real GDP in the United States was $17,688.89 billion in 2016 and $18,108.082 billion in 2017. The total population of the United States was approximately 323.32 million in 2016 and increased to 325.42 million people in 2017.
    1. Calculate the total economic growth rate of the U.S. between 2016 and 2017.
    2. What was the rate of population increase between the two years?
    3. Calculate the levels of Real GDP per capita for the U.S. in 2016 and 2017.
    4. What was the growth rate of Real GDP per capita between 2016 and 2017?
  1. A country currently has a growth rate of Real GDP per capita of 2.6%/year.
    1. According to the Rule of 70, approximately how many years will it take for Real GDP per capita in this economy to double in value?
    2. What growth rate is necessary for this economy to see Real GDP per capita double in only 20 years?
  1. A country has an average population growth rate of 1.5% per year. If Real GDP per capita in this economy is increasing at an annual rate of 3.2% per year, what is the average total economic growth rate for this country?
  1. Differentiate between short-run economic growth and long-run economic growth. According to Neoclassical growth theories, what is the primary determinant of long-run economic growth?

  1. For each of the following institutions, briefly explain how they help to promote economic growth:
    1. Enforceable property rights
    2. Political stability
    3. Public Education

  1. For each situation, indicate the effects on consumption and saving in the economy:
    1. The federal government increases payroll taxes.
    2. The Federal Reserve reduces key interest rate targets.
    3. Consumers are fearful of a possible recession.
    4. There is an increase in rent and mortgage payments for housing.
Answer #1

Total economic growth rate of the U.S. between 2016 and 2017 = (18108.082 - 17688.89)/17688.89

Total economic growth rate of the U.S. between 2016 and 2017 = 2.37%

=======

rate of population increase between the two years = (325.42 - 323.32 )/323.32

rate of population increase between the two years = .65%

=======

Real GDP per capita in 2016 = 17688.89*1000/323.32 = $54710.16

Real GDP per capita in 2017 = 18108.082*1000/325.42 = $55645.26

========

Real GDP per capita growth rate = (55645.26-54710.16)/54710.16

Real GDP per capita growth rate = 1.71%

========

A.

As per the rule of 70,

Number of years required to double  Real GDP per capita = 70/2.6

Number of years required to double  Real GDP per capita = 26.92 years or 27 years

=========

B.

Let, required growth rate = R

Then,

2 = 1*(1+R)^20

R = 2^(1/20) - 1

R = 3.53%

So, with the growth rate of 3.53%, Real GDP per capita doubles in 20 years.

=========

Pl. repost other unanswered questions for their proper answers!